40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 204 of 339  Not logged in ELOG logo
ID Date Author Type Category Subject
  6831   Mon Jun 18 23:38:39 2012 JenneUpdateLSCAdded LSC channels to frames

Since the .ini files get overwritten every time a model is compiled now, we need to put all channels we want saved to frames in the DAQ Channels list inside the model.

I added the _ERR channels for all RFPDs (I and Q for each), as well as the _OUT channels for the DCPDs.  I also added the _OUT channels for the DoF servos (ex. C1:LSC-DARM_OUT).  I don't remember off the top of my head what else we used to save from the LSC model, but those all seemed like ones we'll possibly want access to later. 

We need to go through and do this to all the models we use regularly.

Since SUS hasn't been recompiled in a while, all those channels are saved (until such time as someone does a recompile).  Den has gone through and edited the PEM and OAF .ini files by hand each time he recompiles, so we have that data, although we need to put it into the model (which is the new proper way to acquire channels).

  6830   Mon Jun 18 17:28:03 2012 yutaSummaryComputersbugs in CDS_PARTS/simLinkParts/Fcn

Fcn module in CDS_PARTS is used to include a user defined function in a model.
We should be able to use this by entering desired function, but I found some bugs.

BUG1: Fcn doen't work without ";"

If you put ";" after the function, we can compile.

 sin(u[1]);

But if you put without ";", like

 sin(u[1])

you get the following error message when compiling.

controls@c1ioo
~ 0$ rtcds make c1gcv
### building c1gcv...
Cleaning c1gcv...
Done
Parsing the model c1gcv...
Done
Building EPICS sequencers...
Done
Building front-end Linux kernel module c1gcv...
echo >> target/c1gcvepics/README.making_changes
echo 'Built on date' `date` >> target/c1gcvepics/README.making_changes
make[1]: Leaving directory `/opt/rtcds/caltech/c1/rtbuild'

make[1]: Entering directory `/opt/rtcds/caltech/c1/rtbuild/src/fe/c1gcv'
make -C /lib/modules/2.6.34.1/build SUBDIRS=/opt/rtcds/caltech/c1/rtbuild/src/fe/c1gcv modules
make[2]: Entering directory `/usr/src/linux-2.6.34.1-cs'
  CC [M]  /opt/rtcds/caltech/c1/rtbuild/src/fe/c1gcv/c1gcv.o
make[2]: Leaving directory `/usr/src/linux-2.6.34.1-cs'
make[1]: Leaving directory `/opt/rtcds/caltech/c1/rtbuild/src/fe/c1gcv'
/opt/rtcds/caltech/c1/rtbuild/src/fe/c1gcv/c1gcv.c: In function 'feCode':
/opt/rtcds/caltech/c1/rtbuild/src/fe/c1gcv/c1gcv.c:615: error: expected expression before ';' token
make[3]: *** [/opt/rtcds/caltech/c1/rtbuild/src/fe/c1gcv/c1gcv.o] Error 1
make[2]: *** [_module_/opt/rtcds/caltech/c1/rtbuild/src/fe/c1gcv] Error 2
make[1]: *** [default] Error 2
make: *** [c1gcv] Error 1


BUG2: sindeg doesn't work properly

sindeg should work as cosine with input in degrees.
I made a simple model to test this(below).
model_sindegbug.png


Output of the filter module C1:ALS-BEATY_FINE_PHASE goes to "PHASE_in"
sindeg of this goes to C1:ALS-BEATY_FINE_I_ERR
cosdeg of this goes to C1:ALS-BEATY_FINE_Q_ERR

If you sweep the phase input, you should get sin and cos, but you get the following.
cosdeg (C1:ALS-BEATY_FINE_Q_ERR) looks OK, but sindeg (C1:ALS-BEATY_FINE_I_ERR) looks funny. It looks like ~20000 is its period.

dv_sindegbug.png

  6829   Mon Jun 18 16:23:59 2012 JenneUpdateLockingLSC trigger update

The LSC triggers for the individual filter modules in a filter bank now works.  This is handy so that boosts can come on as soon as a cavity is locked, but will turn off when the cavity unlocks.

You choose which filter modules you want to be triggered, and which ones you want to be manually controlled. 

Example:  LSC-YARM    FM4 and FM5 should always be on, but FM2 and FM3 are controlled by the trigger.  You can set the trigger thresholds for the filter modules independently of the main DoF enable trigger thresholds.

  6828   Mon Jun 18 02:31:43 2012 yutaSummaryGreen Lockinganalysis of mode scan data

I analyzed mode scan data from last week.
Mode matching ratio for Y arm is 86.7 +/- 0.3 %. Assuming we can get rid of TEM01/10 by alignment, this can be improved up to ~ 90%.

Peak search, peak fitting and finnesse calculation:
  I made a python script for doing this. It currently lives in /users/yuta/scripts/modescanresults/analyzemodescan.py.
  What it does is as follows

  1. Read mode scan data(coarse5FSRscan.csv, fine1FSRscan.csv). Each column in the data file should be

[time] [some thing like C1:ALS-BEAT(Y|X)_(COARSE|FINE)_(I|Q)_IN1] [C1:LSC-POY11_I_ERR] [C1:LSC-TRY_OUT]

Each separated by comma. Currently, this script uses only TRY, but it reads all anyway

  2. Find peak in TRY data. For the peak search, it splits data in 1 sec and find local maximum. If the local maximum is higher than given threshold, it recognize it as a peak. If two peaks are very close, it uses higher one. This sometimes fails, because mode scan data we have is not so nice.

  3. Fit each peak with Lorentzian function,

TRY = a*b/(4*(t-c)^2+b^2) + d  (a>0, b>0)

  where a/b is a peak height, b is a linewidth (FWHM), c is a peak position in time, and d is a offset.
  I don't like this, but currently, a/b+c is fixed to the maximum value of TRY data used for fitting. This is because sometimes TRY data is so bad and I couldn't get the peak height correctly. Each points of TRY data doesn't have same error because cavity length is fluctuating and relation between cavity length and TRY is not linear. I think I should use some weighting for the fit, but currently, I just use least squares.

  4. Find TEM00 and calculate FSR in "seconds". I just used "seconds" assuming we did a linear sweep. This script recognize TEM00 from the given threshold.

  5. Calculate finesse using FSR and linewidth of the closest TEM00.

  Below are the result plots from this analysis. Calculated finesse looks quite high (~1000). I think this is from non-linearity in the sweep and error in "measured" line width.
coarse5FSRscan.pngfine1FSRscan.png


Higher order modes and RF sidebands:

  Assuming the curvature of ITMY/ETMY are flat/57.5 m, Y arm length is 38.6 m(FSR 3.9 MHz), positions of HOMs and RF sidebands(11/55 MHz) in frequency domain should look like the plot below.
  The script for calculating this currently lives in /users/yuta/scripts/modescanresults/HOMRFSB.py, inspired by Yoichi's script for KAGRA
HOMRFSB.png

Mode-matching ratio:
  By comparing mode scan data and HOM/RF SB positions in a sophisticated way, you can tell which peak is which.
coarse5FSRscanHOMRFSB.png


  From COARSE 5FSR measurement, peak heights are

TEM00 0.884, 0.896, 0.917, 0.905, 0.911
TEM01 0.040, 0.037, 0.051, 0.054, 0.062
TEM02 0.083, 0.078, 0.079, 0.071, 0.078
TEM03 0.018, 0.015, 0.013, 0.015, 0.014

  So the mode-matching ratio is

MMR = 86.2 %, 87.3 %, 86.5 %, 86.6 %, 85.5 %

  From FINE 1FSR measurement, peak heights and mode matching ratio is

TEM00 0.921
TEM01 0.031
TEM02 0.078
TEM03 0.014

MMR = 88.2 %

  Assuming each measurement had same error, mode-matching ratio from these 6 values is

MMR = 86.7 +/- 0.3 %  (error in 1 sigma)

  This can be improved by ~5% by alignment because we still see ~5% of TEM01/10. Study in systematic errors on going.

  6827   Sat Jun 16 19:32:11 2012 yutaUpdateGreen LockingY arm length using 5FSR scan

I know!
But I think there's some error (~ 10% ?) in calibrating the beatbox. In elog #6815, slope near zero crossing point is about 68 counts/MHz, but now, its 60 counts/MHz. Also, zero crossing point in elog #6815 was 47 MHz, but now, its 45 MHz. 5FSR scan was done between these two calibration measurement.

Quote:

Quote:

Calibrating beat frequency to Y arm length change;
  I used L = 32.36 m (figure above, bottom plot).
    dnu_g / dL = c / lamb_g / L = 1.74 MHz/m

Wow. This is way too short.

You don't need to use Albertoo's arm length as his measurement was done before the upgrade.

 

  6826   Sat Jun 16 18:51:44 2012 KojiUpdateGreen LockingY arm length using 5FSR scan

Quote:

Calibrating beat frequency to Y arm length change;
  I used L = 32.36 m (figure above, bottom plot).
    dnu_g / dL = c / lamb_g / L = 1.74 MHz/m

Wow. This is way too short.

You don't need to use Albertoo's arm length as his measurement was done before the upgrade.

  6825   Sat Jun 16 18:17:00 2012 yutaUpdateGreen LockingY arm length using 5FSR scan

Calibrating error signal to beat frequency;
  I injected 0 dBm RF sine wave into the beatbox and sweeped the frequency(just like we did in elog #6815).
  This time, we have different whitening filters. I sweeped the frequency from 0 to 100 MHz in 200 sec.
  The length of the delay line is ~1.5 m for COARSE.
ALS-BEATY_COARSE_I_IN1_DQ.png

Y arm length;
  Here, I think we need some assumption. Let's assume wavelength of IR lamb_IR = 1064 nm and Y end green frequency is nu_g = 2*nu_IR.
  There is a relation
    dnu_g / nu_g = dL / L
  So,
    dnu_g / (dL/lamb_IR) = 2*nu_IR * lamb_IR / L = 2c/L
  We know that dL/lamb_IR = 1/2 for difference in beat frequency between TEM00s. Therefore, slope of the dnu_g vs dL/lamb_IR plot gives us the arm length L(figure below, middle plot).

CalibYarmScan20120614_2.png

  Error estimation is not done yet, but I think the COARSE_I_IN1 error signal to the beat frequency calibration has the largest error because it seems like the amplitude of sine wave changes ~10% day by day.

Calibrating beat frequency to Y arm length change;
  I used L = 32.36 m (figure above, bottom plot).
    dnu_g / dL = c / lamb_g / L = 1.74 MHz/m

  6824   Sat Jun 16 13:01:17 2012 yutaUpdateGreen Lockingscanned Y arm for 5FSR

Quote:

Is that time stamp really correct?

 Yes. I used pyNDS to get data, but here's a screenshot of dataviewer playing back 300 seconds from GPS time 1023780144.


YarmScanDV.png

  6823   Sat Jun 16 12:03:41 2012 ZachUpdateGreen Lockingscanned Y arm for 5FSR

Is that time stamp really correct? I wanted to look at the signal closely to see if I could get any feeling for why it would look so different when positive vs. negative, but I do not see a triangle anywhere near this time (1023780144)...

Quote:

Interesting. It seems for me that there is a dependence of the noisiness as the beat frequency is scanned.

As you increase (or decrease?) the offset, C1:ALS-BEATY-COARSE_I_IN1 becomes bigger and more crisp.
The resulting out-of-loop stability also seems to be improved as you can see from the crispness of the PDH signal.

Do you find why this happens? Is this because the beat S/N depends on the beat frequency due to the PD noise?

 

 

  6822   Sat Jun 16 01:03:21 2012 yutaUpdateGreen Lockingused longer delay line for mode scan

[Mengyao, Yuta]

Last night, I used 1.5 m delay line COARSE and got 5FSR mode scan. The range 5FSR was limited by the range of SR560.
So, this time, we used 6.4 m(21 feet) cable as a delay line for FINE servo. This should increase the sensitivity by factor of 4. But the range will be 4 tmes smaller, ~ 1.3FSR.

Below is the plot of the mode scan.
You can see the peak height difference between TEM00s, but it's just from the resolution of pixels.

You still can see noisiness goes up when blue plot goes down. But this time, 2000 stands for 27 MHz and -2000 stands for 15 MHz in the beat frequency because we flipped the filter gain this time.
Last night, the top of the triangle was about 40 MHz and bottom was about 60 MHz.


YarmScan20120615.png

We are going to derive mode-matching and some cavity parameters using this plot.

  6821   Fri Jun 15 13:33:39 2012 yutaUpdateGreen LockingADC noise contribution to ALS

ADC noise is not a limiting noise source in a current ALS setup.

Below is the calibrated spectrum of C1:ALS-COARSE_I_ERR when
  Y arm swinging with just damping (red; taken last night)
  terminated before AA (green)
  blocked PSL green beam (blue)

Blue and green curve tells us that noise from the beat PD to ADC is not contributing to the Y arm length sensing noise.

YarmALSnoise20120615.png

  6820   Fri Jun 15 01:53:05 2012 KojiUpdateGreen Lockingscanned Y arm for 5FSR

Interesting. It seems for me that there is a dependence of the noisiness as the beat frequency is scanned.

As you increase (or decrease?) the offset, C1:ALS-BEATY-COARSE_I_IN1 becomes bigger and more crisp.
The resulting out-of-loop stability also seems to be improved as you can see from the crispness of the PDH signal.

Do you find why this happens? Is this because the beat S/N depends on the beat frequency due to the PD noise?

 

  6819   Fri Jun 15 00:50:54 2012 yutaUpdateGreen Lockingscanned Y arm for 5FSR

I scanned Y arm for 5FSR (below).
I could done this after I put a whitening filter.
Currently, whitening filter between the beatbox and AA filter is made of

  Ponoma blue box(passive filter with zero at 1 Hz, pole at 10 Hz) + SR560(flat gain 100)

I couldn't do more than 5FSR because SR560 overloads. I checked it by staring at the indicator during the scan.
Reason why we kept loosing lock last night was the overload of  SR560. Mystery solved!

Anyway, 5FSR is enough.
Our next step is to reduce residual arm length fluctuation.

YarmScan20120614_2.png


Also, I increased the alingnment of IR. So, the higher order modes are less than the last scan.

  6818   Thu Jun 14 21:37:37 2012 yutaUpdateGreen Lockingsucceeded in 1FSR mode scan

[Jenne, Yuta]

We couldn't scan the Y arm for 1FSR last night because the ALS servo breaks while sweeping.
We thought this might be from the amplitude fluctuation of the beat signal. The amplitude of the beat signal goes into the beatbox was about -5 dBm, which is not so enough for the beatbox to get good LO. So, we put an amplifier (and attenuators) and the amplitude became +1 dBm. The range beatbox can handle is about -3 dBm to +3 dBm, according to our calculation.

This increased stability of the lock, and we could scan the arm for 1FSR. Below is the plot of scanned ALS error signal (blue), Y arm IR PDH signal (green) and TRY (red).

YarmScan20120614.png

For each slope, we can see two TEM00 peaks, some higer order modes(may be 01, 02, 02) and sidebands (large 11MHz, small 55MHz?).

We couldn't scan for more. This is still a mystery.

Also, we need to reduce residual Y arm length fluctuation more because we get funny TRY peak shape.

Scan speed:
  For C1:ALS-BEATY_COARSE_I_IN1, 1 count stands for 0.21 nm(see elog #6817). We sweeped 4000 peak to peak in 50 sec. So, the scan speed is about 17 nm/sec.
  This means it takes about 0.06 sec to cross resonant peak.
  Cavity build up time is about 2LF/(pi*c) ~ 40 usec. So, the scan is quasi-static enough.
  Characteristic time scale for the Y end temperature control is about 10 sec, so Y end frequency is following the Y arm length change with temperature control.

  Currently, sampling frequency of DQ channels are 2048 Hz. This means we have 100 points in a TRY peak. I think this is enough to get a peak height.

Next step:
  - Reduce RMS. We are trying to use a whitening filter.
  - Find why we can't scan more. Why??
  - ETMY coil gains may have some unbalance. We need to check
  - Characterize Y end green frequency control. Koji and I changed them last week (see elog #6776).
  - Calculate positions of RF SBs and HOMs and compare with this result.

  6817   Thu Jun 14 04:53:39 2012 yutaSummaryGreen Lockingdesigning ALS loop for mode scan

[[Requirement]]
 Arm cavity FWHM for IR is

  FWHM = FSR / F = c/(2LF) = 8 kHz.

 In cavity length, this is

  L/f * FWHM = 40m/(c/1064nm) = 1.2 nm

 So, to do mode scan nicely, arm length fluctuation during resonant peak crossing should be much less than 1.2 nm.


[[Diagram]]
 Let's consider only ADC noise and seismic noise.
ALSloop.png

* S: conversion from Y arm length to the beat frequency

  dL/L = df/f

 So,

  S = df/dL = f/L = c/532nm/40m = 1.4e7 MHz/m


* W: whitening filter

 We set it to flat gain 50. So,

  W = 50


* D: AD conversion of voltage to counts

 D = 2^16counts/20V = 3300 counts/V


* B: frequency to voltage conversion of the beatbox.

 We measured BWD(elog #6815). When we measured this, W was 10. So, the calibration factor at 0 crossing point(~ 50 MHz) is

  B = 1400*0.048/10/D = 0.0021 V/MHz


* A: actuator transferfunction

 I didn't measure this, but this should look like a simple pendulum with ~ 1 Hz resonant frequency.


* n_ADC: ADC noise

 ADC noise is about

  n_ADC = sqrt(2*LSB^2*Ts) = sqrt(2*(20V/2^14)**2*1/64KHz) = 1.6 uV/rtHz


* n_seis: seismic noise

 We measured this by measuring C1:ALS-BEATY_COARSE_I_IN1. This is actually measuring

  D(WBSn_seis + n_ADC)

 Calibrated plot is the red spectrum below.


* F: servo filter (basically C1:ALS-YARM)

 We need to design this. Stabilized arm length fluctuation is

  x_stab = 1/(1+G)*n_seis + G/(1+G)*n_ADC/(WBS)

 where openloop transferfunction G = SBWDFA.
 Below ~ 50 Hz, n_seis is bigger than n_ADC/(WBS). We don't want to introduce ADC noise to the arm. So, UGF should be around 50 Hz. So, we need phase margin around 50 Hz.
 We also need about 10^3 DC gain to get the first term comparable to the second term.

 Considering these things, openloop transferfunction should look like the below left. Expected error signal when ALS on is the below right. I put some resonant gain to get rid of the peaks which contribute to the RMS (stack at 3.2Hz, bounce at 16.5 Hz).
 Inloop RMS we get is about 0.3 nm, which is only 4 times smaller than FWHM.
ALSopenloop.pngyarmlength.png



[[Discussion]]
 We need to reduce RMS more by factor of ~ 30 to get resolusion 1% of FWHM.
 Most contributing factor to the RMS is power line noise. We might want comb filters, but it's difficult because UGF is at around this region.

 So, I think we need more fancy whitening filters. Currently, we can't increase the gain of the whitening filter because SR560 is almost over loading. Whitening filter with zero at 1 Hz might help.

  6816   Thu Jun 14 01:36:34 2012 yutaUpdateGreen Lockingcan't scan Y arm for 1FSR

[Jenne, Koji, Yuta]

We tried to scan of the Y arm but we couldn't scan for more than 1FSR.
In principle, we can do that because the error signal we are using, C1:ALS-BEATY_COARSE_I_IN1, has the range of ~ 40 MHz, which is about 10FSR (see elog http://nodus.ligo.caltech.edu:8080/40m/6815).

ALS stays for more than 10 min when we don't do the scan. If we put some offset gradually from C1ALS-OFFSETTER2, the lock breaks.
We monitored PZT output of the Y end laser, C1:GCY-SLOW_SERVO1_IN1, but it stayed in the range when scanning. So, there must be something wrong in the ALS loop.

Current in-loop arm length fluctuation is about 0.1 nm RMS (0.5 counts RMS).
Below is the spectrum of the error signal when the ALS is off(green) and on (pink,red). Below ~ 50 Hz, the measurement of the Y arm length is limited by ADC noise (~ 2uV/rtHz).
BEATY_COARSE_LoopOnOff.png

  6815   Wed Jun 13 17:39:13 2012 yutaUpdateGreen Lockingcalibrating the beatbox

[Jenne, Yuta]

We put 0 dBm sine wave to the RF input of the beatbox and linear-sweeped frequency of the sine wave from 0 to 200 MHz using network analyzer (Aligent 4395A).
(We first tried to use 11 MHz EOM marconi)

Whlile the sweep, we recorded the output of the beatbox, C1:ALS-BEATY_(FINE|COARSE)_(I|Q)_IN1_DQ. We made them DQ channels today. Also, we put gain 10 after the beatbox before ADC for temporal whitening filter using SR560s.

We fitted the signals with sine wave using least squares fit(scipy.optimize.leastsq).
Transision time of the frequency from 200 MHz to 0 Hz can be seen from the discontinuity in the time series. We can convert time to frequency using this and supposing linear sweep of the network analyzer is perfect.

Plots below are time series data of each signal(top) and expansion of the fitted region with x axis calibrated in frequency (bottom).

ALS-BEATY_COARSE_I_IN1_DQ.pngALS-BEATY_COARSE_Q_IN1_DQ.png
ALS-BEATY_FINE_I_IN1_DQ.pngALS-BEATY_FINE_Q_IN1_DQ.png


We got

C1:ALS-BEATY_COARSE_I_IN1_DQ = -1400 sin(0.048 freq + 1.17pi) - 410
C1:ALS-BEATY_COARSE_Q_IN1_DQ = 1900 sin(0.045 freq + 0.80pi) - 95

C1:ALS-BEATY_FINE_I_IN1_DQ = 1400 sin(0.89 freq + 0.74pi) + 15
C1:ALS-BEATY_FINE_Q_IN1_DQ = 1400 sin(0.89 freq + 1.24pi) - 3.4

(freq in MHz)

The delay line length calculated from this fitted value (supposing speed of signal in cable is 0.7c) is;

  D_coarse = 0.7c * 0.048/(2*pi*1MHz) =  1.6 m
  D_fine = 0.7c * 0.89/(2*pi*1MHz) = 30 m

So, the measurement look quite reasonable.

FINE signals looks nice because we have similar response with 0.5pi phase difference.
For COARSE, maybe we need to do the measurement again because the frequency discontinuity may affected the shape of the signal.

  6814   Wed Jun 13 11:19:05 2012 steveUpdateSUSIDC receptacles clamped

The MC_ IDC 64 pin cables from sat. amplifiers to junction-interface-board towards  whitening - dewhitening at the back of rack 1 X 5 are finally  clamped with

All other sus cables of the same kind have the correct short latch arm to lock them in for reliable contact.

Attachment 1: IMG_1341.JPG
IMG_1341.JPG
  6813   Wed Jun 13 11:10:56 2012 ranaUpdateGreen Lockingmy first modescan (sort of)

You can easily calculate whether or not the coarse readout will work by thinking about the scan resolution you need given the ADC dynamic range and the whitening filter that you use.

  6812   Wed Jun 13 03:03:38 2012 yutaUpdateGreen Lockingmy first modescan (sort of)

Linear range df of the delay line technique is about df ~ c/(2D). So, the linear range for the fine signal(delay line length D=30m) is about 5 MHz.
Arm cavity FSR = c/(2L) = 3.7 MHz.
So, I think we need phase shifting to do mode scan for more than 2 FSRs by holding the arm length finely with fine servo.
For the coarse (D=1.5m), the linear range is about 100 MHz, so if we can do mode scan using coarse servo, it is OK.

In any case, I think it is nice to have linear signal with fixed slope even if we don't adjust the phase every time.

Quote:

 That sounds goofy.

With the delay line technique, you can get a linear signal over 50 MHz with no phase shifting. What is with all this I/Q stuff?

 

  6811   Wed Jun 13 02:24:02 2012 ranaUpdateGreen Lockingmy first modescan (sort of)

 That sounds goofy.

With the delay line technique, you can get a linear signal over 50 MHz with no phase shifting. What is with all this I/Q stuff?

  6810   Wed Jun 13 02:11:59 2012 yutaUpdateGreen Lockingmy first modescan (sort of)

I stabilized Y arm length by using only I phase coarse signal from the beat(C1:ALS-BEATY_COARSE_I_ERR).
I sweeped the arm length by injecting 0.05Hz sine wave from C1:ALS_OFFSETTER2_EXC.
Below is the plot of TRY and the error signal(ideally, Y arm length) while the sweep.

modescan20120612_1.png

I couldn't hold the arm length tight, so you can see multiple peaks close to each other.
We need to
  - adjust offsets
  - adjust rotation phase of I-Q mixing
  - adjust servo filters

to hold the length tighter.

Also, I couldn't sweep the Y arm length very much. I need to calibrate, but to do the modescan for many FSRs, we need to
  - introduce automatic phase optimizing system
There were sin/cos function in the CDS_PARTS, so I think we can feedback I_ERR to control rotation phase of I-Q mixing.

  6809   Tue Jun 12 23:18:18 2012 yutaUpdateGreen LockingI-Q signals for the beat

[Mengyao, Yuta]

Yes!! We have I-Q signals for the beat!!

What we did:
  1. Aligned Y arm to the Y end green incident beam. The transmission to the PSL was about 195 uW.

  2. Aligned IR beam to the Y arm by adjusting PZTs and got the transmission, C1:LSC-TRY_OUT ~ 0.86.

  3. Aligned green optics on the PSL table to get the beat signal. The beat was found when;

  PSL laser temperature on display: 31.41 deg C
  C1:PSL-FSS_SLOWDC = 1.43
  Y end laser "T+": 34.049 deg C
  Y end laser "ADJ": 0
  Y end laser measured temperature: 34.14 deg C
  C1:GCY-SLOW_SERVO2_OFFSET = 29950
  Y end slow servo: off (was on)

  4. Connected the beat PD output to the beatbox.

  5. Kicked ETMY position to change the cavity length and while the ringdown, we run pynds to get data. We plotted C1:ALS-BEATY_FINE_I_ERR vs C1:ALS-BEATY_FINE_Q_ERR, and C1:ALS-BEATY_COARSE_I_ERR vs C1:ALS-BEATY_COARSE_Q_ERR (below). We got nice circle as expected.

FINEIQplot20120612.pngCOARSEIQplot20120612.png

Current setup:
  Only AA filers are put between the output of the beatbox and the ADC.

beatysetup20120612.png

  6808   Tue Jun 12 20:35:46 2012 yutaUpdateGreen Lockingc1gcv recompiled

[Jamie, Yuta]

We recompiled c1gcv because the order of the channels were confusing. We found some change in the phase rotation module when we did this.

I did some cabling and checked each signals are actually going to the right channel. I labeled all the cables I know, which go into the AA chasis for ADC1 of c1ioo machine.

Below is the list of the channels. If you know anything about "unknown" channels, please let me know.

Current channel assignments for ADC1 of c1ioo machine:
  Red ones were added today. Green ones existed in the past, but channel assignment were changed.

cable

# on AA chassis name in Simulink channel name

connected
but unknown

J1A    
   
not connected J1B    
   
not connected J2 adc_1_2 C1:ALS-XARM_BEAT_DC
not connected adc_1_3 C1:ALS-YARM_BEAT_DC
connected
but unknown
J3    
   
connected
but unknown
J4    
   
connected
but unknown
J5    
   
connected
but unknown
J6    
   
connected
but unknown
J7    
   
beat Y arm fine I J8A adc_1_14 C1:ALS-BEATY_FINE_I
beat Y arm fine Q adc_1_15 C1:ALS-BEATY_FINE_Q
not connected J8B    
   
connected
but unknown
J9A    
   
not connected J9B    
   
connected
but unknown
J10    
   
connected
but unknown
J11    
   
not connected J12 adc_1_22 C1:ALS-BEATX_COARSE_I
not connected adc_1_23 C1:ALS-BEATX_COARSE_Q
not connected J13 adc_1_24 C1:ALS-BEATX_FINE_I
not connected adc_1_25 C1:ALS-BEATX_FINE_Q
beat Y arm coarse I
J14 adc_1_26 C1:ALS-BEATY_COARSE_I
beat Y arm coarse Q adc_1_27 C1:ALS-BEATY_COARSE_Q
not connected J15 adc_1_28 Broken! Don't use this!!
adc_1_29 (not broken)
not connected J16A adc_1_30 (not broken)
adc_1_31 Broken? Funny signal.
not connected J16B    
   

Memorandum for me:
  Recompiling procedure;

ssh c1ioo

rtcds make c1gcv
rtcds install c1gcv
rtcds start c1gcv

Attachment 1: c1gcv20120612-2.png
c1gcv20120612-2.png
  6807   Tue Jun 12 17:46:09 2012 JenneUpdateComputersrtcds: command found

Quote:

Quote:

We can't compile any changes to the LSC or the GCV models since Jamie's new script / program isn't found.  I don't know where it is (I can't find it either), so I can't do the compiling by hand, or point explicitly to the script.  The old way of compiling models in the wiki is obsolete, and didn't work :(

Sorry about that.  I had modified the path environment that pointed to the rtcds util.  The rtcds util is now in /opt/rtcds/caltech/c1/scripts/rtcds, which is in the path.  Starting a new shell should make it available again.

 Added TRX and TRY and POY11_I_ERR and POX11_I_ERR to the c1lsc.mdl using a new-style DAQ Channels block, recompiled, installed, started the model, all good.  Restarted the daqd on the framebuilder, and everything is green.  I can go back and get recorded data using dataviewer (for the last few minutes since I started fb), so it all looks good.

Note on the new DAQ Channels block:  Put the text block (from CDS_PARTS) at the same level as the channel you want to save, and name it exactly as it is in the model.  The code-generator will add the _DQ for you.  i.e. if you define a channel "TRY_OUT_DQ" in the lsc model, you'll end up with a channel "C1:LSC-TRY_OUT_DQ_DQ".

  6806   Tue Jun 12 17:29:28 2012 DenUpdateCDSdq channels

All PEM and IOO DQ channels disappeared. These channels were commented in C1???.ini files though I've uncommented them a few weeks ago. It happened after these models were rebuild, C1???.ini files also changed. Why?

I added the channels back. mx_stream died on c1sus after I pressed DAQ reload on medm screen. For IOO model it is even worse. After pressing DAQ Reload for C1IOO model DACQ process dies on the FB and IOO machine suspends.

I rebooted IOO, restarted models and fb. Models work now, but there might be an easier way to add channels without rebooting machines and demons.

  6805   Tue Jun 12 17:04:55 2012 steveUpdateSUSPRM oplevs servo is still bad

Quote:

Quote:

Yuta claims he fixed the PRM oplev by centering it the other day, but no one has left it on and watched it for a long while, to make sure it's okay.  We watched it now for ~2 min, and it was good, but we're leaving the oplevs off anyway for the night.  Tomorrow we should restore PRM (it's currently restored), turn on the oplevs, and let it sit to make sure it doesn't go crazy.

 

 PRM oplev servo was turned on with PITgain 0.5  and YAWgain  -0.7

Note: gain settings were PIT  1.0  and  YAW --0.5   on Jun 1, 2012 that I measured Feb 23, 2012

 It is still oscillating. Gains turned down to zero.

Attachment 1: PRMolvstillosc.png
PRMolvstillosc.png
  6804   Tue Jun 12 16:33:32 2012 steveUpdateSUSPRM oplevs servo ON for confirmation

Quote:

Yuta claims he fixed the PRM oplev by centering it the other day, but no one has left it on and watched it for a long while, to make sure it's okay.  We watched it now for ~2 min, and it was good, but we're leaving the oplevs off anyway for the night.  Tomorrow we should restore PRM (it's currently restored), turn on the oplevs, and let it sit to make sure it doesn't go crazy.

 

 PRM oplev servo was turned on with PITgain 0.5  and YAWgain  -0.7

Note: gain settings were PIT  1.0  and  YAW --0.5   on Jun 1, 2012 that I measured Feb 23, 2012

  6803   Tue Jun 12 13:49:32 2012 JamieConfigurationComputer Scripts / Programstconvert

A nicer, better maintained version of tconvert is now supplied by the lalapps package.  It's called lalapps_tconvert.  I installed lalapps on all the workstations and aliased tconvert to point to lalapps_tconvert.

  6802   Tue Jun 12 11:54:50 2012 JenneUpdateGreen Lockingc1gcv recompiled

Yuta added channels so we can get the Q phase of all the beat PDs to the c1gcv model.  I showed him how to recompile/install/start.

During the install, it couldn't find: Unable to find the following file in CDS_MEDM_PATH: LOCKIN_FILTER.adl

On all the screens (ALS and SUS), lockin parts are white.  Someone changed something, then didn't go back to fix the screens.

Otherwise, things look to be working fine.

Attachment 1: c1gcv20120612.png
c1gcv20120612.png
  6801   Tue Jun 12 11:25:05 2012 JamieUpdateComputersrtcds: command not found

Quote:

We can't compile any changes to the LSC or the GCV models since Jamie's new script / program isn't found.  I don't know where it is (I can't find it either), so I can't do the compiling by hand, or point explicitly to the script.  The old way of compiling models in the wiki is obsolete, and didn't work :(

Sorry about that.  I had modified the path environment that pointed to the rtcds util.  The rtcds util is now in /opt/rtcds/caltech/c1/scripts/rtcds, which is in the path.  Starting a new shell should make it available again.

  6800   Tue Jun 12 02:09:43 2012 JenneUpdateComputersmini boot fest

Quote:

stock-vector-wine-icon-on-computer-keyboard-original-illustration-58731781.jpg

 I'm starting to feel like a wine-o here.  Yuta wanted to glance at the PRM oplev dataviewer, and lo and behold, the fb lost connection just as he decided to do that.  We had checked the front end status screen not 1 minute beforehand, and everything was green.  Lame.

  6799   Tue Jun 12 02:07:42 2012 JenneUpdateSUSPRM oplevs left off

Yuta claims he fixed the PRM oplev by centering it the other day, but no one has left it on and watched it for a long while, to make sure it's okay.  We watched it now for ~2 min, and it was good, but we're leaving the oplevs off anyway for the night.  Tomorrow we should restore PRM (it's currently restored), turn on the oplevs, and let it sit to make sure it doesn't go crazy.

 

  6798   Tue Jun 12 01:58:33 2012 yutaUpdateGreen Lockingaligned Y arm to Y end green

[Jenne, Yuta]

We aligned Y arm to the Y end green incident beam.
We noticed two TEM00, bright and dim, so we decreased Y end laser temperature to 34.13 deg C.
It doubled the transmission of the green, and now the transmission to the PSL table is 178 uW, which is close to the maximum(197 uW) we got so far.

Current settings for Y end laser is;

  Y end laser "T+": 34.049 deg C
  Y end laser "ADJ": 0
  Y end laser measured temperature: 34.13 deg C
  C1:GCY-SLOW_SERVO2_OFFSET = 31025
  Y end slow servo: on (was off)

We aligned IR beam to the Y arm by mostly adjusting PZTs and got the transmission, C1:LSC-TRY_OUT ~ 0.9.

We tried to calculate the mode-matching ratio for IR by taking TRY data while ITMY and ETMY are swinging (without ALS), but it was difficult because we see too many higher order modes.

Tomorrow, we will (1) connect the beatbox to ADC, (2) edit c1gcv model, (3) scan the arm using I-Q signals.

  6797   Tue Jun 12 01:03:18 2012 JenneUpdateComputersmini boot fest

As usual, we noticed the frame builder wasn't connecting happily with the rest of the computers just as we were about to lock some stuff (we never notice it being bad when we're not trying to use the frame builder....)

All the big rectangles by each computer were white.  I restarted daqd, and that brought most things back.  c1lsc and c1sus needed their mx_streams restarted manually to get everything green again.

stock-vector-wine-icon-on-computer-keyboard-original-illustration-58731781.jpg

I've had about enough whine with my computers for tonight.

  6796   Tue Jun 12 00:15:51 2012 JenneUpdateComputersrtcds: command not found

We can't compile any changes to the LSC or the GCV models since Jamie's new script / program isn't found.  I don't know where it is (I can't find it either), so I can't do the compiling by hand, or point explicitly to the script.  The old way of compiling models in the wiki is obsolete, and didn't work :(

This means we can't (a) record TRY or (b) add the Q quadrature of the beat PD to the real time system tonight.

We're going to try just using Yuta's pynds script to capture data in real time, so we can keep working for tonight.

  6795   Mon Jun 11 22:51:11 2012 JenneUpdateComputerstdsavg not working

tdsavg isn't working:

controls@rossa:/opt/rtcds/caltech/c1/scripts/LSC 6$ tdsavg 10 C1:LSC-ASDC_IN1
ERROR: LDAQ - Unable to find NDS host "fb0"
ERROR: LDAQ - Unable to find NDS host "fb1"
ERROR: LDAQ - Unable to open socket to NDS.

 

When this command is executed inside a script, it doesn't return anything.  eg:

set offset = `tdsavg 10 C1:LSC_ASDC_IN1`
echo $offset

returns a blank line. 

Past elog research said lots of things about test points.  I didn't suspect that, since there aren't many test points occupied (according to the CDS status screens), but I cleared the test points anyway (elog 6319).  Didn't change anything, still broken.

LSCoffsets script, and any others depending on tdsavg will not work until this is fixed.

  6794   Mon Jun 11 21:50:08 2012 yutaUpdateGreen Lockingbeatbox looks OK

Summary:
  We need I-Q frequency deiscriminator to control the arm length fine and continuously.
  I checked the beatbox (LIGO-D1102241-v4; see elog #6302) and it was working.

What I did:
  1. Measured some transferfunctions with a network analyzer (Aligent 4395A) and checked the cabling is correct.

  2. Put 30 m/1.5 m delay line and checked I-Q outputs are actually orthogonal. I did this by sweeping the frequency of RF input to the beatbox. See attached picture. You can see nice circle on the oscilloscope.

Some measurement results:

  - Gains of the transferfunctions(@ 10-100MHz) between;

   RF in -> RF mon: -25 to -20 dB
   RF in -> fine delay out: -50 to -40 dB
   RF in -> coarse delay out: -50 to -40 dB
   RF in -> LO of mixer RMS-1: ~ +4 dB  (RMS-1 needs +7 dB LO)
 
  - 30m delay line(RG-142B/U) had -2 dB loss.

Note:
  - RF input must be larger than about -3 dBm to get enough LO to the mixer. Otherwise, you won't get I-Q outputs.
  - The comparator, whitening filter and differential DAQ outputs are not installed in the current beatbox.
  - Current beatbox only has electronics for the one arm.
  - The print on the board D1102241 says +15V and -15V, but they are actually opposite. Cabling is swapped in order to supply correct power to the ICs.

Attachment 1: CIMG1522.JPG
CIMG1522.JPG
  6793   Mon Jun 11 21:35:55 2012 JenneUpdateEnvironmentRattling in the HEPA

Quote:

There is an intermittent rattling sound coming from the HEPA in the NE corner of the PSL table (right above the PMC, all of our input optics).

Steve says it might be a bad bearing, but he'll check it out in the morning and get it fixed.

 MC was having a hard time staying locked, with no discernable reason from the control room (i.e. no big seismic, no PMC PZT railing).    The HEPA was on 100%, so I turned it down to 50% to hopefully reduce the rattling, if that was what was wrong. 

  6792   Mon Jun 11 16:08:58 2012 JenneUpdateEnvironmentRattling in the HEPA

There is an intermittent rattling sound coming from the HEPA in the NE corner of the PSL table (right above the PMC, all of our input optics).

Steve says it might be a bad bearing, but he'll check it out in the morning and get it fixed.

  6791   Mon Jun 11 09:37:16 2012 steveUpdateIOOPMC locked

Quote:

Quote:

IOO Angle & IOO Position QPDs centered.

 PMC trend of 400 and 1200 days

The Innolight 2W based PSL- IOO was implemented in the ~ summer of 2010

 The PMC was locked and the MC followed intantly

Attachment 1: PMClocked.png
PMClocked.png
  6790   Fri Jun 8 15:24:08 2012 steveUpdateVACRGA scan at day 276 and pressure plots

Quote:

We tried not to open chambers above 10,000 particles of 0.5 micron cf/min

New items going in:                       2 rasor beam traps, 5 badly oxidized old silver plated  setscrew with spring loaded tips......to be replaced in the future, viton tips for eq screws....some are lose, gold plated allen wrench installed at ITMX bottom, reglued magnet on ITMX

Bad hardware things found:          nylon ball "locking elements" on OSEM locking set screws with screwdriver  slot, lose  1064 nm filter on OSEM pd

 40m - pumpdown 71- maglev pumping speed - at day 276........all normal

Attachment 1: 40m71pdM276d.png
40m71pdM276d.png
Attachment 2: pd71m276d.png
pd71m276d.png
Attachment 3: pd71d276.png
pd71d276.png
  6789   Fri Jun 8 15:08:27 2012 yutaUpdateGreen Lockingaligned/mode-matched Y green beat setup

Laser temperature settings for Y arm green work today are;

  PSL laser temperature on display: 31.38 deg C (PSL HEPA 100%)
  C1:PSL-FSS_SLOWDC = 1.68
  Y end laser "T+": 34.049 deg C
  Y end laser "ADJ": 0
  Y end laser measured temperature: 34.13 deg C (*)
  C1:GCY-SLOW_SERVO2_OFFSET = 29845

Green transmission from Y end and PSL green power on the beat PD are;

  P_Y = 28 uW
  P_PSL = 96 uW

P_Y decrease from its maximum we got (75 uW, see elog #6777) is because the alignment for Y arm green is decreased. I can see the decrease from the green reflection on ETMT camera, but I will leave it because we already have enough beat.

I aligned PSL optics, including the mode-matching lens to maximize the beat note. The beat note I got is about 26dBm.
The calculated value is -14 dBm, so we have about 75 % loss.
I measured the reflection from the PD window and its reflectivity was about 30%. We still have unknown 45% loss.

  6788   Thu Jun 7 18:46:13 2012 yutaUpdateSUSPRM oplev centered

PRM oplev beam was not hitting on the QPD since Jun 1, so I centered it. I reverted the oplev servo gains and now oplev servo looks fine.

C1:SUS-PRM_OLPIT_GAIN = 1.0
C1:SUS-PRM_OLYAW_GAIN = -0.7

There's SIDE to UL/UR/LL/LR coil element in PRM TO_COIL matrix. They were 0 until Mar 31, 2012, but someone changed them to -0.160. I couldn't find elog about it.
Same thing happened to BS on Mar 13, 2012 (see elog #6409), so I think somebody did the same thing to PRM.

  6787   Thu Jun 7 17:49:09 2012 JamieUpdateCDSc1sus in weird state, running models but unresponsive otherwise

Somehow c1sus was in a very strange state.  It was running models, but EPICS was slow to respond.  We could not log into it via ssh, and we could not bring up test points.  Since we didn't know what else to do we just gave it a hard reset.

Once it came it, none of the models were running.  I think this is a separate problem with the model startup scripts that I need to debug.  I logged on to c1sus and ran:

rtcds restart all

(which handles proper order of restarts) and everything came up fine.

Have no idea what happened there to make c1sus freeze like that.  Will keep an eye out.

  6786   Thu Jun 7 13:25:37 2012 KojiUpdateSUSPRM is still oscillating

Done.

C1:SUS-PRM_OLPIT_GAIN 1.0 -> 0
C1:SUS-PRM_OLYAW_GAIN -0.7 -> 0
 

Quote:

  Set the PRM OL servo gains to zero until someone can take care of this. Turning off the buttons doesn't help anything if people run the alignment scripts.

 

  6785   Thu Jun 7 13:18:42 2012 ranaUpdateSUSPRM is still oscillating

  Set the PRM OL servo gains to zero until someone can take care of this. Turning off the buttons doesn't help anything if people run the alignment scripts.

  6784   Thu Jun 7 12:49:50 2012 JamieUpdateComputerszita network configured

Added zita to the martian host table, and configured his network accordingly:

zita            A       192.168.113.217

https://wiki-40m.ligo.caltech.edu/Martian_Host_Table

  6783   Thu Jun 7 10:39:21 2012 JamieUpdateComputerspianosa dual monitor working

I decided to remove what I thought was the problematic extra nVidia video card, since there are already two DVI outputs build-in.  The card turned out to not even be nVidia, so I don't know what was going on there.

I futzed with the BIOS to configure the primary video card, which is some new Intel card.  The lucid (10.04) support for it is lacking, but it was easy enough to pull in new drivers from the appropriate Ubuntu PPA repository:

controls@pianosa:~ 0$ sudo apt-add-repository ppa:f-hackenberger/x220-intel-mesa
controls@pianosa:~ 0$ sudo apt-add-repository ppa:glasen/intel-driver
controls@pianosa:~ 0$ sudo apt-get update
...
controls@pianosa:~ 0$ sudo apt-get upgrade
Reading package lists... Done
Building dependency tree       
Reading state information... Done
The following packages will be upgraded:
  libdrm-intel1 libdrm-nouveau1 libdrm-radeon1 libdrm2 libgl1-mesa-dri libgl1-mesa-glx libglu1-mesa xserver-xorg-video-intel
8 upgraded, 0 newly installed, 0 to remove and 0 not upgraded.
Need to get 3,212kB of archives.
After this operation, 25.2MB disk space will be freed.
Do you want to continue [Y/n]?
...
controls@pianosa:~ 0$

After a reboot, both monitors came up fine.

http://www.subcritical.org/running_ubuntu_lts_on_sandy_bridge/

  6782   Thu Jun 7 09:52:05 2012 steveUpdatePEMair cond maintenance

Quote:

 

 Air conditioning maintenance is scheduled for tomorrow from 8 to 11am

 Jeff checked and  replaced filters  as needed. Job completed this morning.

Attachment 1: PEM800d.png
PEM800d.png
ELOG V3.1.3-